OpenStudy (anonymous):

Will give medal and become fan (accepting only correct answers)! Solve for x. (x^4 - 1) / (x^3) = 0

OpenStudy (anonymous):

Please explain how you got the answer.

OpenStudy (zpupster):

multply both sides by x^3 factor (x^2-1)(x^2+1) and again (x^2-1) = (x+1)(x-1) set to zero x+1=0 x-1=0 What are your roots??? (x^2+1) can never be 0 when y=0 x is?? see graph

Nnesha (nnesha):

okay so move x^3 to the right side as ~zpupster said multiply both sides by x^3

OpenStudy (anonymous):

Okay so then it would be \[x ^{3} \times (x ^{4} - 1) \div x ^{3} = 0\]

OpenStudy (anonymous):

But the x^3 can be cancelled so... the answer would be:\[x ^{4} - 1 = 0\]

OpenStudy (anonymous):

Right?

Nnesha (nnesha):

yep you need to multiply *both sides* so 0 times x^3 = 0 now apply difference of square rule

Nnesha (nnesha):

\[\huge\rm a^2-b^2 =(a+b)(a-b)\]

OpenStudy (anonymous):

\[(x ^{2} + 1)(x ^{2} - 1)\]

OpenStudy (anonymous):

\[(x ^{2}+1)(x+1)(x-1)\]

OpenStudy (anonymous):

Is that it?

Nnesha (nnesha):

yep right you would apply difference when there is negative sign so x^2-1 apply difference of square

Nnesha (nnesha):

oopps you already know

Nnesha (nnesha):

yes right set all 3 parentheses equal to zero solve for x

OpenStudy (anonymous):

How do you find the square root of -1?

Nnesha (nnesha):

well do you need for this question ?

Nnesha (nnesha):

ohh gotcha so you will get an imaginary solution

OpenStudy (anonymous):

Is it going to be \[\pm 1\]

Nnesha (nnesha):

\[\huge rm \sqrt{-1} = i\]

OpenStudy (anonymous):

Oh like that! So the answer is x = 1, -1, and i

Nnesha (nnesha):

\[\sqrt{-1}=i \]

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Alright thanks again!

Nnesha (nnesha):

my pleasure :=)